OpenStudy (fran345):

Given the function, f(x) = 2x^2 - x + 1, calculate the average rate of change of the function from x = 1 to x = 3. I need help!

OpenStudy (jackthegreatest):

@fran345 So I am assuming the change in y

OpenStudy (jackthegreatest):

@fran345 in which you would first plug x in 2(1)^2 - 1 + 1 = 4 2(3)^2 - 3 + 1 = 34

OpenStudy (jackthegreatest):

changes from 1 to 3 = 30 30/2 = 15

OpenStudy (fran345):

Is your answer 15? I don't have that for my choices.